# Exploring Graph Representations of Logical Forms for Language Modeling

### v1

@@ Abstract
We make the case for language models over logical forms (LFLMs), arguing that such models are more data-efficient than their textual counterparts. To that end, we introduce the Graph-based Formal-Logical Distributional Semantics (GFoLDS) prototype, a pretrained LM over graph representations of logical forms, as a proof-of-concept of LFLMs. Using GFoLDS, we present strong experimental evidence that LFLMs can leverage the built-in, basic linguistic knowledge inherent in such models to immediately begin learning more complex patterns. On downstream tasks, we show that GFoLDS vastly outperforms textual, transformer LMs pretrained on similar amounts of data, indicating that LFLMs can learn with substantially less data than models over plain text. Furthermore, we show that the performance of this model is likely to scale with additional parameters and pretraining data, suggesting the viability of LFLMs in real-world applications.
@@ 1 Introduction
However, there is a considerable amount of evidence in the literature (e.g. Xu et al., 2021; Wu et al., 2021b; Prange et al., 2022; Sachan et al., 2021; Zhou et al., 2020; Zhang et al., 2020, 2022, etc.) indicating that linguistically-informed LMs—models whose inputs and/or architectures are augmented by linguistic knowledge—can improve performance without consuming more (text) data. In this paper, we argue for the use of LMs over logical forms (LFLMs): LMs that take as input semantic representations, rather than text. In particular, we posit the following hypothesis:
@@
The (aspects of) linguistic knowledge incorporated into LFLMs obviates the need to learn elementary linguistic phenomena, allowing them to immediately begin learning more complex patterns.
One relevant corollary of the LKCH is that LFLMs can learn with less data: the basic linguistic knowledge built into LFLMs facilitates more rapid learning of advanced phenomena.
To make the case for LFLMs, we introduce the Graph-based Formal-Logical Distributional Semantics (GFoLDS; Section 3) model [^1] All code used in this paper available on GitHub: https://github.com/mjs227/GFoLDS as a proof-of-concept of such LMs. GFoLDS is a pretrained, encoder graph transformer Wu et al. (2021a) over structures derived from Dependency Minimal Recursion Semantics (DMRS; Copestake, 2009) representations (see Figure 1): directed, labeled acyclic graphs in which nodes correspond to predicates and edge labels denote relations (i.e. semantic roles) between them, in a similar manner to Abstract Meaning Representation (AMR; Banarescu et al., 2013) structures. Unlike AMR, which abstracts away from semantic features such as tense and number, DMRS includes these features, resulting in a more faithful representation of linguistic meaning.
@@
Secondly, we demonstrate the viability of pretrained LFLMs in Section 5, by comparing the performance of GFoLDS to that of BERT (trained on $\sim$6.5 times more data than GFoLDS; Devlin et al., 2019) on a range of downstream tasks. Although the actual BERT models outperform GFoLDS, our model outperforms—by a wide margin—BERT models pretrained on the same data as GFoLDS, indicating that LFLMs can learn useful representations with much less data than their textual counterparts.
@@ 2 Related Work
There exists a body of research indicating that injecting graph representations of linguistic structures into textual LMs can improve downstream performance. Xu et al. (2021) fuse dependency parse graphs into pretrainined transformer encoders (e.g. BERT and RoBERTa; Liu et al., 2019), and surpass the then-SoTA results on relation classification, entity typing, and question answering tasks, demonstrating the general utility of linguistically-informed LMs. Similarly, Wu et al. (2021b) inject syntactic dependency parse graphs into a pretrained BERT model, yielding then-SoTA results on semantic role labeling and relation extraction tasks. These authors additionally compare the respective impacts of syntactic and semantic representations on performance, and find that semantic representations are more beneficial with respect to downstream performance than syntactic structures.
@@
To the best of our knowledge, Functional Distributional Semantics at Scale (FDSAS; Lo et al., 2023) represents the only [^2] Aside from its predecessor: Functional Distributional Semantics Emerson (2018). extant LFLM aside from GFoLDS. FDSAS is a variational autoencoder over DMRS graphs that learns probabilistic truth-conditional functions for each predicate (node label). FDSAS is not, however, a transformer and so does not benefit from the flexibility afforded by such architectures: it is unclear how to scale the model to yield a deeper architecture, and it does not generate usable hidden states that can be passed to (for example) a classification head for downstream tasks. Additionally, its rigorous formal-semantic foundations give rise to a degree of inflexibility in this model: Lo et al. (2023) discarded prepositions, quantifiers, and modal verbs from FDSAS’ training dataset out of necessity.
@@ 3 GFoLDS
In this section, we describe the GFoLDS model architecture (Section 3.1), preprocessing steps that we employed on the DMRS graphs (Section 3.2), and the model’s pretraining procedure (Section 3.3).
@@ 3.1 Architecture
The GFoLDS model is a variant of the graph transformer paradigm Wu et al. (2021a), which was originally introduced for molecule graph classification. A graph transformer consists of a graph neural network (GNN) that encodes local neighborhood information, whose output is then fed to a permutation-invariant (i.e. without linear positional embeddings) transformer encoder for global message-passing (attention). Unique to this work is the GNN component of GFoLDS, which consists of an embedding layer and the positional encoding network (see Figure 2). The output of the embedding layer is fed into the positional encoding network, which provides each node with a representation of its local neighborhood in the DMRS graph structure.
@@
In their respective adaptations of the GraphSAGE (Hamilton et al., 2017) and Graph Convolutional Network (GCN; Kipf and Welling, 2017) architectures to directed graphs, Xu et al. (2018) and Tong et al. (2020) introduce forward and backward node projection layers, which encode information about incoming and outgoing connections (respectively) for a given node. In a similar fashion, each SWA layer contains a forward (Equation 2) and a backward SWA block (see Figure 3), which encode the nodes—and the semantic roles thereof—mapping into and out of a given node.
@@
The backward SWA block is architecturally identical to the forward block, but contains distinct edge projection layers $W_{\ell}^{(b)}$ and operates on the transpose $G^{T}$ of the graph. The respective forward and backward representations of each node $n_{i}$ are summed together and passed through a two layer, feed-forward block (identical to the feed-forward blocks in the model’s encoder layers).
@@ 3.2 Data Preprocessing
GFoLDS’ pretraining corpus consisted of $\sim$17.5 million randomly-selected sentences from the November 1, 2023 English Wikipedia dump [^3] https://huggingface.co/datasets/wikimedia/wikipedia; CC-BY-SA-3.0 license., constituting a total of $\sim$508 million words ($\sim$6.5 times smaller than BERT’s pre-training corpus). We first used Spacy’s SentenceRecognizer [^4] https://spacy.io/api/sentencerecognizer pipeline to extract individual sentences from the text. We then used the PyDelphin (Goodman, 2019) library with the ACE/ERG (Copestake and Flickinger, 2000) rule-based parser/grammar [^5] ERG-1214 release: https://github.com/delph-in/erg to obtain a DMRS representation of each sentence, before preprocessing the resulting DMRS structures to yield GFoLDS input graphs. The ACE/ERG parser was able to parse $\sim$84of the data, for a total of $\sim$14.6 million DMRS-derived graphs.
It was not feasible to tokenize named entities (CARGs) and out-of-vocabulary (OOV) items in the same manner as for the in-vocabulary DMRS predicates—i.e. by simply assigning an integer to each unique predicate string in the vocabulary. We therefore replaced all CARGS and OOV terms with the [MASK] token. These [MASK] tokens are not targets for prediction during the pre-training procedure (as they are OOV, so there is no possible target token); the goal is instead to have the model represent the OOV item with the closest in-vocabulary token, based on the context in which the OOV item appears. The removal of CARGs and OOV items is a stop-gap measure, and remains an open problem and barrier to the performance of the GFoLDS model: we defer the incorporation of CARGs and OOV items into the model’s input structures to future work (see Section 8.1).
@@ 4 Evaluating the LKCH
This section is dedicated to an investigation of the validity of the LKCH. The hypothesis can be broken down into two distinct claims: (i) that the (aspects of) linguistic knowledge incorporated into LFLMs obviates the need to learn elementary linguistic phenomena; and (ii) that this enables such models to then immediately begin learning more complex patterns.
We therefore divided this experiment into two parts, which respectively probe the model’s knowledge of elementary and complex linguistic phenomena. To validate the claim made in the LKCH that linguistically-informed LMs immediately begin to learn complex patterns, we evaluated the model at regular intervals throughout pretraining.
@@ 4.2 Experimental Setup
This experiment consisted of two tasks that were designed to evaluate knowledge of elementary linguistic phenomena, and one to evaluate knowledge of more complex patterns. We evaluated GFoLDS and the comparison models on the three tasks at twenty evenly-spaced intervals per pretraining epoch, for a total of eighty points of comparison.
Assuming that the LKCH holds, we should expect to see GFoLDS outperform the BERT-C models on the complex task at each stage of the pre-training process, as—according to the hypothesis—GFoLDS is able to learn complex patterns faster than textual LMs. On the elementary tasks, we expect GFoLDS to outperform BERT-C, but also that GFoLDS’s performance will not improve significantly across pretraining, while BERT-C’s performance steadily increases: the LKCH asserts that an LFLM is already equipped with elementary linguistic knowledge, so its performance at the first checkpoint should be near (or above) that of a textual model such as BERT-C at the last checkpoint.
@@ 4.2.1 Complex Task
The complex task in this experiment is the RELPRON Rimell et al. (2016) dataset. This dataset consists of terms (nouns), each paired with a hypernym and up to ten properties: relative clauses that restrict that hypernym (see Table 2 in the Appendix). The development set consists of 65 terms and 518 properties ($\sim$8 properties per term on average), and the test set contains 73 terms and 569 properties ($\sim$7.8 per term). The task is to retrieve the properties that apply to each term, without including those that do not: the evaluation metric is Mean Average Precision (MAP) score. Due to the small size of the dataset (and the lack of a training split), the frozen, pre-trained models were used to obtain token probabilities for property ranking (see Appendix D for additional details).
To evaluate GFoLDS, we discarded all examples containing CARG-bearing predicates or OOV items (the “RELPRON-No-UNK/NE” row of Table 1). As discussed in Section 3.2, CARGs and OOV items are masked in the input graphs: given that each template only contains four content words (the hypernym, verb, relative clause subject/object, and the target term), the GFoLDS model is effectively blind to (at least) one third of the context in templates that contain OOV items or CARGs.
After discarding all examples that contain CARGs or OOV items, the resulting No-UNK/NE subset of the RELPRON test split contains 63 terms and 421 properties, for a total of $\sim$6.68 properties per term on average.
Figure 4: Precision scores on the two elementary tasks (left and center) and MAP scores on the RELPRON test set (right) across the 80 evenly-spaced training snapshots for GFoLDS and the BERT comparison models.
@@ 4.2.2 Elementary Tasks
POS-Prediction
The first task, POS-prediction, evaluates the LMs’ ability to model the distribution of parts-of-speech: this is a commonly-employed probing task used to assess LMs’ elementary linguistic competence (Waldis et al., 2024). We evaluated the models on 200 sentences (not drawn from their pretraining data) from English Wikipedia. For each sentence, we masked a single word and recorded its POS: we then extracted the models’ probability distribution over the masked word of each sentence, and recorded their precision at ten with respect to the masked word’s part-of-speech. For example, if the masked word was a noun, and nine of the model’s top ten most-likely predictions were nouns, then the model received a score of 0.9 for that instance.
Quantifier-Agreement
The second elementary task, quantifier-agreement, evaluates the models’ knowledge of quantifier number agreement—whether a quantifier restricts a singular or plural noun (or both)—across 179 sentences drawn from English Wikipedia (not used to pretrain the models). Although less common than POS prediction, quantifier agreement tasks have also been employed as a metric of language models’ elementary lingustic competence (see e.g. Huebner et al., 2021; Waldis et al., 2024).
@@
Additional details on the setup and evaluation of the elementary tasks are given in Appendix G.
@@ 4.3 Results
These results constitute strong evidence towards the LKCH. It is clear from this experiment that GFoLDS can model the elementary phenomena from the onset (and retains this ability throughout pretraining) and GFoLDS’ performance on the RELPRON test set suggests that this translates to more rapid learning of more complex patterns.

### v2

@@ Abstract
We make the case for language models over logical forms (LFLMs), arguing that such models are more data-efficient than their textual counterparts. To that end, we introduce the Graph-based Formal-Logical Distributional Semantics (GFoLDS) prototype, a pretrained LM over graph representations of logical forms, as a proof-of-concept of LFLMs. Using GFoLDS, we present strong experimental evidence that LFLMs can leverage the built-in, basic linguistic knowledge inherent in such models to immediately begin learning more complex patterns. On downstream tasks, we show that GFoLDS vastly outperforms textual, transformer LMs (BERT) pretrained on the same data, indicating that LFLMs can learn with substantially less data than models over plain text. Furthermore, we show that the performance of this model is likely to scale with additional parameters and pretraining data, suggesting the viability of LFLMs in real-world applications.
@@ 1 Introduction
However, there is a considerable amount of evidence in the literature (e.g. Xu et al., 2021; Wu et al., 2021b; Prange et al., 2022; Sachan et al., 2021; Zhou et al., 2020; Zhang et al., 2020, 2022, etc.) indicating that linguistically-informed LMs—models whose inputs and/or architectures are augmented by linguistic knowledge—can improve performance without consuming more text data.
In this paper, we argue for the use of LMs over logical forms (LFLMs): LMs that take as input semantic representations, rather than text. In particular, we posit the following hypothesis:
@@
The (aspects of) linguistic knowledge incorporated into LFLMs greatly accelerates their learning of elementary linguistic phenomena, in turn accelerating the learning of more complex patterns.
One relevant corollary of the LKCH is that LFLMs can learn with less data: the linguistic knowledge built into LFLMs facilitates more rapid learning of advanced phenomena.
We argue that the primary advantage of logical forms (as opposed to other types of linguistic knowledge) with respect to language modeling is the de-noising effect conferred by the function-argument structure inherent to such representations. Specifically, the translation of surface text to logical form has an equivalence-classing effect, so that all syntactic paraphrases of the same proposition—for example, an active sentence and its passive counterpart—are mapped to the same representation. The benefit to an LM of the de-noising effect resulting from this equivalence-classing is clear: the model does not need to learn to equate periphrastic structures, and so can immediately begin learning co-occurrence relations between predicates.
To make the case empirically for LFLMs, we introduce the Graph-based Formal-Logical Distributional Semantics (GFoLDS; Section 3) model as a proof-of-concept of such LMs. GFoLDS is a pretrained, encoder graph transformer Wu et al. (2021a) over structures derived from Dependency Minimal Recursion Semantics (DMRS; Copestake, 2009) representations (see Figure 1): directed, labeled acyclic graphs in which nodes correspond to predicates and edge labels denote relations (i.e. semantic roles) between them, in a similar manner to Abstract Meaning Representation (AMR; Banarescu et al., 2013) structures.
[…]
@@
Secondly, we demonstrate the viability of pretrained LFLMs in Section 5, by comparing the performance of GFoLDS to that of BERT (trained on $\sim$6.5 times more data than GFoLDS; Devlin et al., 2019) on a range of downstream tasks. Although the actual BERT models outperform GFoLDS, our model outperforms—by a wide margin—BERT models pretrained on the same data as GFoLDS on all benchmarks, indicating that LFLMs can learn useful representations with much less data than their textual counterparts.
@@
We make all code for the GFoLDS model and the experiments conducted in this paper available on GitHub [^1] https://github.com/mjs227/GFoLDS.
@@ 2 Related Work
As discussed in Section 1, there exists a body of research indicating that injecting graph representations of linguistic structures into textual LMs can improve downstream performance. Xu et al. (2021) fuse dependency parse graphs into pretrainined transformer encoders (e.g. BERT and RoBERTa; Liu et al., 2019), and surpass the then-SoTA results on relation classification, entity typing, and question answering tasks, demonstrating the general utility of linguistically-informed LMs. Similarly, Wu et al. (2021b) inject syntactic dependency parse graphs into a pretrained BERT model, yielding then-SoTA results on semantic role labeling and relation extraction tasks. These authors additionally compare the respective impacts of syntactic and semantic representations on performance, and find that semantic representations are more beneficial with respect to downstream performance than syntactic structures.
@@
To the best of our knowledge, Functional Distributional Semantics at Scale (FDSAS; Lo et al., 2023) represents the only [^2] Aside from its predecessor: Functional Distributional Semantics Emerson (2018). extant LFLM aside from GFoLDS. FDSAS is a variational autoencoder over DMRS graphs that learns probabilistic truth-conditional functions for each predicate (node label). FDSAS is not, however, a transformer and so does not benefit from the flexibility afforded by such architectures: it is unclear how to scale the model to yield a deeper architecture, and it does not generate usable hidden states that can be passed to (for example) a classification head for downstream tasks. Additionally, its rigorous formal-semantic foundations give rise to a degree of inflexibility in this model: for example, Lo et al. (2023) discarded prepositions, quantifiers, and modal verbs from FDSAS’ training dataset out of necessity.
@@ 3 GFoLDS
In this section, we describe the GFoLDS model architecture (Section 3.1), preprocessing steps that we employed on its DMRS graph inputs (Section 3.2), and the model’s pretraining procedure (Section 3.3).
@@ 3.1 Architecture
The GFoLDS model is a variant of the graph transformer paradigm Wu et al. (2021a), which was originally introduced for molecule graph classification. A graph transformer consists of a graph neural network (GNN) that encodes local neighborhood information, whose output is then fed to a permutation-invariant (i.e. without linear positional embeddings) transformer encoder for global message-passing (attention).
Unique to this work is the GNN component of GFoLDS, which consists of an embedding layer and the positional encoding network (see Figure 2). The output of the embedding layer is fed into the positional encoding network, which provides each node with a representation of its local neighborhood in the DMRS graph structure.
@@
In their respective adaptations of the GraphSAGE (Hamilton et al., 2017) and Graph Convolutional Network (GCN; Kipf and Welling, 2017) architectures to directed graphs, Xu et al. (2018) and Tong et al. (2020) introduce forward and backward node projection layers, which encode information about incoming and outgoing connections (respectively) for a given node. In a similar fashion, each SWA layer (see Figure 3) contains a forward (Equation 2) and a backward SWA block, which encode the nodes—and the semantic roles thereof—mapping into and out of a given node.
@@
The backward SWA block is architecturally identical to the forward block, but contains distinct edge projection matrices $W_{\ell}^{(b)}$ and operates on the transpose $G^{T}$ of the graph. The respective forward and backward representations of each node $n_{i}$ are summed together and passed through a two layer, feed-forward block (identical to the feed-forward blocks in the model’s encoder layers).
@@ 3.2 Data Preprocessing
GFoLDS’ pretraining corpus consisted of $\sim$17.5 million randomly-selected sentences from the November 1, 2023 English Wikipedia dump [^3] https://huggingface.co/datasets/wikimedia/wikipedia; CC-BY-SA-3.0 license., constituting a total of $\sim$508 million words ($\sim$6.5 times smaller than BERT’s pretraining corpus). We first used Spacy’s SentenceRecognizer [^4] https://spacy.io/api/sentencerecognizer pipeline to extract individual sentences from the text. We then used the PyDelphin (Goodman, 2019) library with the ACE/ERG (Copestake and Flickinger, 2000) rule-based parser/grammar [^5] ERG-1214 release: https://github.com/delph-in/erg to obtain a DMRS representation of each sentence, before preprocessing the resulting DMRS structures to yield GFoLDS input graphs. The ACE/ERG parser was able to parse $\sim$84of the data, for a total of $\sim$14.6 million DMRS-derived graphs.
It was not feasible to tokenize named entities (CARGs) and out-of-vocabulary (OOV) items in the same manner as for the in-vocabulary DMRS predicates—i.e. by simply assigning an integer to each unique predicate string in the vocabulary. We therefore replaced all CARGs and OOV terms with the [MASK] token. These [MASK] tokens are not targets for prediction during the pretraining procedure (as they are OOV, so there is no possible target token): the goal is instead to have the model represent the OOV item with the closest in-vocabulary token, based on the context in which the OOV item appears. The removal of CARGs and OOV items is a stop-gap measure, and remains an open problem and barrier to the performance of the GFoLDS model. We defer the incorporation of CARGs and OOV items into the model’s input structures to future work (see Section 8.1).
@@ 4 Evaluating the LKCH
This section is dedicated to an investigation of the validity of the LKCH. The hypothesis can be broken down into two distinct claims: (i) that the (aspects of) linguistic knowledge incorporated into LFLMs greatly accelerates their learning of elementary linguistic phenomena; and (ii) that this in turn accelerates the learning of more complex patterns.
We therefore divided this experiment into two parts, which respectively probe the model’s knowledge of elementary and complex linguistic phenomena. Due to the claim made in the LKCH that linguistically-informed LMs’ learning of complex patterns is accelerated, we evaluate the model at regular intervals throughout pretraining, in order to measure the rate at which it is learning.
@@ 4.2 Experimental Setup
This experiment consisted of two tasks that were designed to evaluate knowledge of elementary linguistic phenomena, and one to evaluate knowledge of more complex patterns: we evaluated GFoLDS and the BERT comparison models on the three tasks at twenty evenly-spaced intervals per pretraining epoch, for a total of eighty points of comparison.
Note that it is difficult—if not impossible—to define “elementary” and “complex” linguistic phenomena in absolute terms. In this paper, we consider these terms in a relative sense: the elementary tasks (Section 4.2.2) are undoubtedly more elementary than the complex task (Section 4.2.1), in that they do not require as much (if any) world knowledge, and deal entirely with awareness of basic linguistic categories. Conversely, this entails that our complex task is complex relative to our elementary tasks.
Assuming that the LKCH holds, we should expect to see GFoLDS outperform the BERT-C models on the complex task throughout the pretraining process, as—according to the hypothesis—GFoLDS is able to learn complex patterns faster than textual LMs.
On the elementary tasks, we again expect GFoLDS to outperform BERT-C, but also that GFoLDS’ performance will improve substantially faster than it does on the complex tasks: the LKCH predicts that an LFLM’s accelerated learning of elementary phenomena catalyzes its learning of complex patterns, so its learning of the former should therefore accelerate at a faster rate than that of the latter.
@@ 4.2.1 Complex Task
The complex task in this experiment is the RELPRON Rimell et al. (2016) dataset. This dataset consists of terms (nouns), each paired with a hypernym and up to ten properties: relative clauses that restrict that hypernym (see Table 2 in the Appendix). The development set consists of 65 terms and 518 properties ($\sim$8 properties per term on average), and the test set contains 73 terms and 569 properties ($\sim$7.8 per term). The task is to retrieve the properties that apply to each term, without including those that do not: the evaluation metric is Mean Average Precision (MAP) score.
To evaluate the models, we constructed templates out of each (term, hypernym, property) triple: for example, the triple (telescope, device, astronomers use) yields the template “a device that astronomers use is a telescope”. We then replaced the target term with the [MASK] token (e.g. “a device that astronomers use is a [MASK]”): the probability assigned to a given term under the masked distribution is taken as proportional to the probability that the property applies to that term.
Figure 4: Precision scores on the two elementary tasks (left and center) and MAP scores on the RELPRON test set (right) across the 80 evenly-spaced training snapshots for GFoLDS (green) and the $\text{BERT}_{\text{base}}$/$\text{BERT}_{\text{large}}$ comparison models (blue/orange, respectively).
As discussed in Section 3.2, CARGs and OOV items are masked in the input graphs: given that each template only contains four content words (the hypernym, verb, relative clause subject/object, and the target term), the GFoLDS model is effectively blind to (at least) one third of the context in templates that contain OOV items or CARGs. We therefore discarded all examples containing CARG-bearing predicates or OOV items for evaluation. This resulting subset of the test split (the “RELPRON-No-UNK/NE” column of Table 1) contains 63 terms and 421 properties, for a total of $\sim$6.68 properties per term on average.
Due to the small size of the dataset (and the lack of a training split), the frozen, pretrained models were used to obtain token probabilities for property ranking on this task.
@@ 4.2.2 Elementary Tasks
POS-Prediction.
The first elementary task evaluates the LMs’ ability to model the distribution of parts-of-speech: this is a commonly-employed probing task used to assess LMs’ elementary linguistic competence (Waldis et al., 2024). We evaluated the models on 200 sentences (not drawn from their pretraining data) from English Wikipedia. For each sentence, we masked a single word and recorded its POS: we then extracted the models’ probability distribution over the masked word of each sentence, and recorded their precision at ten with respect to the masked word’s part-of-speech. For example, if the masked word was a noun, and nine of the model’s top ten most-likely predictions were nouns, then the model received a score of 0.9 for that instance.
Quantifier-Agreement.
The second elementary task evaluates the models’ knowledge of quantifier number agreement—whether a quantifier restricts a singular or plural noun (or both)—across 179 sentences drawn from English Wikipedia (not used to pretrain the models). Although less common than POS prediction, quantifier agreement tasks have also been employed as a metric of language models’ elementary lingustic competence (see e.g. Huebner et al., 2021; Waldis et al., 2024).
@@
Additional details on the setup and evaluation of the elementary tasks are given in Appendix F.
@@ 4.3 Results
The results of this experiment (Figure 4) conform almost exactly to the behavior predicted by the LKCH: the performance of GFoLDS on the elementary tasks remains relatively constant throughout pretraining, because the model starts near peak performance from the first checkpoint—this indicates that its learning of elementary patterns was complete within 5% of the first epoch. While the performance of the BERT models steadily increases on these tasks, it does so at a much lower rate.
@@
These results constitute strong evidence towards the LKCH. It is clear from this experiment that GFoLDS can model the elementary phenomena almost from the onset (and retains this ability throughout pretraining) and GFoLDS’ performance on the RELPRON test set suggests that this accelerated learning of elementary phenomena translates to more rapid learning of more complex patterns.